OpenStudy (anonymous):

Given the quadratic function f(x) = x^2 – 6x + 9, find a value of x such that f(x) = 25

OpenStudy (lalaly):

\[x^2-6x+9=25\]\[x^2-6x-16=0\] \[(x-8)(x+2)=0\]

OpenStudy (lalaly):

x=8 and x=-2

OpenStudy (anonymous):

f(x)= x^2 – 6x + 9 = 0 => f(25) = 625 -6(25)+9 = 484

OpenStudy (anonymous):

so this is the right answer

OpenStudy (lalaly):

raheen f(x)=25 not f(25)

OpenStudy (anonymous):

oh. you are right. I keep facing lag in this website